A bus starts from rest and accelerates at 1.5 m/s2 until it reaches a velocity of 9m/s. The bus continues at this velocity and then accelerates at -2 m/s2 until it comes to a stop 400m from its starting point. How much time did the bus take to cover the 400m?


1
Expert's answer
2021-05-09T13:03:25-0400

"s_1=v^2/(2a_1)=9^2/(2\cdot1.5)=27\ (m)" ; "v=a_1t_1\to t_1=v/a_1=9/1.5=6\ (s)"


"s_3=v^2/(2a_3)=9^2/(2\cdot2)=20.25\ (m)" ; "v'=v-a_3t_3\to t_3=v/a_3=9/2=4.5\ (s)"


"s_2=400-27-20.25=352.75\ (m)" ; "t_2=s_2/v=352.75/9=39.19\ (s)"


"t=6+4.5+39.19=49.69\ (s)" . Answer